# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v3

@@ 5.1 Experimental Settings
Baselines. We compare with state-of-the-art prompt optimization methods as follows:
  • GPT [2] series are powerful large language models capable of zero-shot prompt optimization. We use GPT-4o-mini for lower cost and better efficiency.
  • MagicPrompt [17] trains GPT-2 on high-quality prompts collected from stable-diffusion model users.
  • Promptist [26] optimizes user input into model-preferred prompts via reinforcement learning.
@@
Evaluation Metrics. We employ four latest metrics FDD [61], Aesthetic Score [19], AI Corrupt Score [45], and Vendi Score [22] to measure the quality of generated images. Specifically, FDD (Frechet Dino Distance) measures the image fidelity using the similarity between generated and dataset images via DinoV2 features [47], which better aligns with human perception than traditional FID [28]. Aesthetic Score is computed via Aesthetic Predictor V2.5 [19], quantifying visual appeal, composition quality, and artistic merit. AI Corrupt Score detects technical flaws in generated images by identifying visual artifacts. Vendi Score quantifies image diversity by calculating the von Neumann entropy from a normalized cosine similarity matrix using DinoV2 embeddings.
@@
Evaluation Protocols. Our proposed TIPO leverages large-scale image caption datasets for training, which overlap with the training text distributions of many T2I models. Following Promptist [26], we divide our experiments into two settings: (1) In-domain, where the T2I model’s training texts overlap with those used by TIPO, and (2) Out-of-domain, where no overlap exists.
5.2 In-domain Tag-based Prompt Optimization
To assess prompt optimization performance on tag-based prompts, we generate scenery images, as they contain abundant descriptive tags for objects and backgrounds. We randomly sample 32,768 tag-based prompts from Danbooru2023 [69, 46], shuffle and concatenate the scenery tags into new prompts (thereby preventing data leakage of the original captions), and generate one image per prompt using Kohaku-XL-Zeta, an SDXL [50, 46] variant fine-tuned on Danbooru2023. This evaluation tests the in-domain capabilities, as Danbooru2023 is also used during TIPO training.
Metric Original GPT MagicPrompt Promptist TIPO (Ours)
FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
Aesthetic Score $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
AI Corrupt Score $\uparrow$ 0.4257 0.7490 0.5024 0.5669 0.9195
Vendi Score $\uparrow$ 16.8135 8.6632 11.9014 14.3273 13.3070
Table 1: Performance on in-domain tag-based prompts, with best results bolded and the second-best underlined.
The results in Table 1 reveal two key insights. First, MagicPrompt and Promptist, which rely on user prompts or reinforcement learning, underperform in Aesthetic and AI Corrupted Scores due to the quality or quantity limitations of their collected samples (e.g., Promptist uses 90K samples, limited by the high reinforcement learning cost). In contrast, GPT and TIPO benefit from large-scale training corpora ($>$30M samples), yielding higher-quality outputs. Second, TIPO achieves the best FDD by a substantial margin over GPT, which can be attributed to its superior distribution alignment with T2I models.
5.3 In-domain NL-based Prompt Optimization
We evaluate the prompt optimization performance on NL-based prompts by selecting 10,000 short prompts and 10,000 long prompts from CaptionEmporium [12] and GBC [29] as test prompts. In particular, since the long prompts are much longer than typical user input, we truncate them to two sentences ($<$ 40 words) to simulate real-world applications. We use SDXL-1.0-base [50] as the T2I model, whose training text data largely overlap with TIPO.
Short Original GPT MagicPrompt Promptist TIPO(Ours)
FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
Aesthetic Score $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
AI Corrupt Score $\uparrow$ 0.7113 0.6985 0.7064 0.6314 0.7131
Vendi Score $\uparrow$ 38.1715 34.7139 38.1553 34.1270 37.0649
Truncated Long Original GPT MagicPrompt Promptist TIPO(Ours)
FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
Aesthetic Score $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
AI Corrupt Score $\uparrow$ 0.6868 0.6712 0.6741 0.5925 0.7130
Vendi Score $\uparrow$ 38.2533 34.8108 37.8412 33.5266 37.0900
Table 2: Performance on in-domain NL-based prompts.
@@ 5.3 In-domain NL-based Prompt Optimization
Figure 5: Generated images from 4 types of prompts: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated ($<$ 40 words) long prompt, (d) TIPO-enhanced truncated prompt. TIPO adds detail and maintains variety, yielding coherent images from simple prompts.
@@
Table 2 demonstrates that TIPO achieves either the best or second-best scores in Aesthetic and AI Corrupt Score by effectively enriching the original prompt with appropriate textual elements while rarely introducing extraneous noise. While all methods compromise fidelity and diversity, as reflected in the FDD and Vendi Score, TIPO remains competitive because it maintains small semantic deviation from the original sentences via progressive refinement.
(a) Oringial Prompt
(b) Text Completion
(c) TIPO
Figure 6: Qualitative comparison of prompt completion versus TIPO (ours). (a) shows images generated by original prompts. (b) depicts images generated by text completion using TIPO-500M. (c) shows images generated by TIPO.
5.4 Out-of-domain Performance
@@
Some recent T2I models are trained on proprietary images and captions. As a representative example, SD-3.5-Large[20] is trained on private images captioned with CogView[71], which differ markedly from the texts used to train TIPO. To evaluate model performance in this out-of-domain scenario, we generate 8,192 original tag- and NL-based prompts using the baseline GPT-4o-mini rather than relying on existing prompt datasets. We apply the remaining methods to these prompts and assess their performance.
GPT MagicPrompt Promptist TIPO(Ours)
Aesthetic Score $\uparrow$ 6.7125 6.4507 6.3924 6.0536
AI Corrupt Score $\uparrow$ 0.9482 0.8577 0.9053 0.9280
Vendi Score $\uparrow$ 8.9718 15.8721 16.4891 21.5706
Table 3: Performance on the out-of-domain T2I model
@@
5.5 Human Preference Evaluation
@@
Quantitative metrics may not fully align with human preference. Therefore, we conducted a user study based on pairwise image comparisons between the original prompt, MagicPrompt, Promptist, and TIPO on over 1,400 images, gathering preferences from 221 volunteers. As illustrated in Figure 7, TIPO achieved the highest overall win rate at 51.3%, significantly outperforming competitors. In out-of-domain scenarios, TIPO’s win rate increased to 52.5%, demonstrating consistently strong user preference across different contexts. For further results and statistics, please refer to Appendix F.
Figure 7: The win-tie-lose rate on overall preference on all data and out-of-domain situation where images are generated by SD-3.5-Medium.
Figure 8: ELO rating on overall preference on all data
5.6 Ablation Study
Original Completion TIPO
Aesthetic Score $\uparrow$ 5.9507 5.9631 5.9960
Corrupt Score $\uparrow$ 0.6799 0.8131 0.8971
Vendi Score $\uparrow$ 19.7208 18.2896 13.9864
Table 4: Ablation results
We attribute TIPO’s success primarily to its multi-task pre-sampling design detailed in Section 4.3. However, a natural question arises: can a simple text completion approach achieve similar or better results? To validate our design, we compare TIPO’s pre-sampling strategy with a direct text completion using the same language model, TIPO-500M. As shown in Table 4, TIPO significantly improves aesthetic scores and reduces artifacts compared to text completion. For visual examples, please refer to Figure 6. For more ablation details and visualization, please refer to Appendix 5.6.
@@ 6 Conclusion
We introduced TIPO, a lightweight prompt pre-sampling framework designed for efficient real-world Text-to-Image (T2I) applications. By aligning user prompts with the intrinsic distributions of T2I training datasets, TIPO enhances semantic coherence, image fidelity, and diversity with minimal inference overhead.
TIPO consistently outperforms existing prompt optimization methods, achieving superior performance across multiple evaluation metrics. Its ability to preserve user intent while enhancing prompt diversity and image quality makes it well-suited for both in-domain and out-of-domain scenarios, demonstrating robustness across various prompt formats and model architectures.
Extensive user studies also confirmed TIPO’s strong alignment with human preferences, further highlighting its potential for practical applications. To encourage wider adoption and facilitate reproducibility, we release our trained models and source code. We hope TIPO will inspire further advancements in efficient, scalable, and robust generative frameworks for creative systems.
@@ A.1 Danbooru2023
The Danbooru2023 dataset [69, 68, 46] is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board. This dataset includes images annotated with particular and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework. The dataset contains data up to image ID 7,349,999, encompassing various visual content with granular annotations. These annotations allow for creating nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.

### v4

@@ 5.1 Experimental Settings
Baselines and T2I Models
We compare against representative prompt-optimization baselines: GPT-4o-mini for zero-shot rewriting (OpenAI, 2024), MagicPrompt, which fine-tunes GPT-2 on community-collected prompts (daspartho, 2022), Promptist, which applies reinforcement learning for model-preferred prompt optimization (Hao et al., 2023), and Gemini-2.0-flash-image  [^3] https://developers.googleblog.com/experiment-with-gemini-20-flash-native-image-generation/, which uniquely serves both as a prompt-refinement baseline and a T2I generator. For image generation backbones, our main experiments adopt SDXL-base-1.0 (Podell et al., 2024), Illustrious v3.5 (v-pred), Kohaku-XL-Zeta, and Stable Diffusion 3.5 Large(Esser et al., 2024a). To further assess generalization, we additionally evaluate on four diverse backbones with undisclosed training data: FLUX.1-dev (black-forest-labs, 2024), Omnigen2 (Wu et al., 2025), Lumina-2 (Qin et al., 2025), and HiDream-I1 (Cai et al., 2025). The details of T2I models are provided in Appendix B.
Evaluation Metrics
@@
We employ four latest metrics FDD (Stein et al., 2023), Aesthetic Score (discus0434, 2024), AI Corrupt Score (narugo1992, 2023), and Vendi Score (Friedman and Dieng, 2022) to measure the quality of generated images. Specifically, FDD (Fréchet DINO Distance) quantifies fidelity by comparing the distribution of DINOv2 features (Oquab et al., 2023) between reference images in the evaluation dataset and images generated from the corresponding captions, which better aligns with human perception than traditional FID (Heusel et al., 2017). Aesthetic Score is computed via Aesthetic Predictor V2.5 (discus0434, 2024), quantifying visual appeal, composition quality, and artistic merit. AI Corrupt Score detects technical flaws in generated images by identifying visual artifacts. Vendi Score quantifies image diversity by calculating the von Neumann entropy from a normalized cosine similarity matrix using DinoV2 embeddings. Notably, Vendi is defined directly on feature-space dispersion and is sensitive to non-semantic variations such as low-level noise or artifacts. As a result, it “should be used alongside a quality metric” (Friedman and Dieng, 2022) and a higher value alone does not always correspond to more meaningful semantic diversity.
Evaluation Protocols
@@
Our proposed TIPO leverages large-scale image caption datasets for training, which overlap with the training text distributions of many T2I models. Following Promptist (Hao et al., 2023), we divide our experiments into two settings: (1) In-domain, where the T2I model’s training texts overlap with those used by TIPO, and (2) Out-of-domain, where no overlap exists.
@@
Figure 4: Generated images from 4 types of prompts: (a) simple scenery tag, (b) scenery tag enhanced by TIPO, (c) truncated (<< 40 words) long prompt, (d) TIPO-enhanced truncated prompt. TIPO adds detail and maintains variety, yielding coherent images from simple prompts.
5.2 Experimental Results
In-domain Tag-based Prompt Optimization
To assess prompt optimization performance on tag-based prompts, we generate scenery images, as they contain abundant descriptive tags for objects and backgrounds. We randomly sample 32,768 tag-based prompts from Danbooru2023 (Yeh, 2024a; nyanko202, 2023), shuffle and concatenate the scenery tags into new prompts (thereby preventing data leakage of the original captions), and generate one image per prompt using Kohaku-XL-Zeta. This evaluation tests the in-domain capabilities, as Danbooru2023 is used during both TIPO and Kohaku-XL-Zeta training. The results in Table 2 reveal two key insights. First, MagicPrompt and Promptist, which rely on user prompts or reinforcement learning, underperform in Aesthetic and AI Corrupted Scores due to the quality or quantity limitations of their collected samples (e.g., Promptist uses 90K samples, limited by the high reinforcement learning cost). In contrast, GPT and TIPO benefit from large-scale training corpora (>>30M samples), yielding higher-quality outputs. Second, TIPO achieves the best FDD by a substantial margin over GPT, which can be attributed to its superior distribution alignment with T2I models.
Table 2: Comprehensive performance comparison of TIPO against baselines across different in-domain prompt types. Metrics are marked with $\uparrow$ (higher is better) or $\downarrow$ (lower is better). For OOD tests, the ’Original’ baseline and the FDD metric was not applicable. In the table, Aesthetic refers to Aesthetic Score, Corrupt to AI Corrupt Score, and Vendi to Vendi Score. TIPO demonstrates significant improvements, achieving the highest average rank among all baselines. Best results are in bold and second-best are underlined.
Prompt Type / Task Metric Original GPT MagicPrompt Promptist TIPO
In-domain Tag-based Prompts FDD $\downarrow$ 0.3558 0.5414 0.3247 0.2350 0.2282
Aesthetic $\uparrow$ 5.0569 6.3676 6.1609 5.9468 6.2571
Corrupt $\downarrow$ 0.5743 0.2510 0.4976 0.4331 0.0805
Vendi $\uparrow$ 16.814 8.663 11.901 14.327 13.307
In-domain NL-based (Short) FDD $\downarrow$ 0.0957 0.1668 0.0980 0.1783 0.1168
Aesthetic $\uparrow$ 5.8370 6.0589 5.8213 5.7963 5.8531
Corrupt $\downarrow$ 0.2887 0.3015 0.2936 0.3686 0.2870
Vendi $\uparrow$ 38.172 34.714 38.155 34.127 37.065
In-domain NL-based (Truncated Long) FDD $\downarrow$ 0.0955 0.1683 0.1247 0.2096 0.1210
Aesthetic $\uparrow$ 5.7497 6.0168 5.8191 5.7759 5.8364
Corrupt $\downarrow$ 0.3132 0.3288 0.3259 0.4075 0.2870
Vendi $\uparrow$ 38.253 34.811 37.841 33.527 37.090
Out-of-Domain (OOD) Test Aesthetic $\uparrow$ N/A 6.7125 6.4507 6.3924 6.0536
Corrupt $\downarrow$ N/A 0.0518 0.1423 0.0947 0.0720
Vendi $\uparrow$ N/A 8.9718 15.872 16.489 21.571
Overall Average Rank $\downarrow$ 2.58 3.00 3.00 3.87 2.07
Table 3: TIPO on T2I models with undisclosed training data. Significant improvements are in bold.
Model Variant Aesthetic ↑ AI Corrupt ↓ FDD ↓ Vendi ↑
FLUX.1-dev Original 5.2029 0.1202 0.1185 36.2597
TIPO 5.2746 0.0938 0.1202 35.6489
Omnigen2 Original 5.2629 0.1110 0.1373 33.7724
TIPO 5.0661 0.1187 0.1253 34.2681
Lumina-2 Original 5.2588 0.0981 0.1318 34.7599
TIPO 5.4105 0.0916 0.1286 34.4655
HiDream-I1 Original 5.7768 0.1055 0.1444 34.6318
TIPO 5.8143 0.1019 0.1379 34.3544
Gemini-2.0-Flash-Image Self-refined 5.2584 0.0928 0.8084 32.7422
TIPO 5.2649 0.0700 0.7964 32.1136
In-domain NL-based Prompt Optimization
@@ 5.2 Experimental Results
We evaluate the prompt optimization performance on NL-based prompts by selecting 10,000 short prompts and 10,000 long prompts from CaptionEmporium (CaptionEmporium, 2024) and GBC (Hsieh et al., 2024) as test prompts. In particular, since the long prompts are much longer than typical user input, we truncate them to two sentences (<< 40 words) to simulate real-world applications. We use SDXL-1.0-base as the T2I model, whose training text data largely overlap with TIPO. Table 2 demonstrates that TIPO achieves either the best or second-best scores in Aesthetic and AI Corrupt Score by effectively enriching the original prompt with appropriate textual elements while rarely introducing extraneous noise. While all methods compromise fidelity and diversity, as reflected in the FDD and Vendi Score, TIPO remains competitive because it maintains small semantic deviation from the original sentences via progressive refinement.
Out-of-domain Performance
@@
Some recent T2I models are trained on proprietary images and captions. As a representative example, SD-3.5-Large is trained on private images captioned with CogView (Zheng et al., 2024), which differ markedly from the texts used to train TIPO. To evaluate model performance in this out-of-domain scenario, we generate 8,192 original tag- and NL-based prompts using the baseline GPT-4o-mini rather than relying on existing prompt datasets. We apply the remaining methods to these prompts and assess their performance.
@@
Compatibility with other T2I Models
In addition to the above baselines with publicly known training data, we further evaluate TIPO on recent models with undisclosed training sources: FLUX.1-dev (black-forest-labs, 2024), Omnigen2 (Wu et al., 2025), Lumina-2 (Qin et al., 2025), HiDream-I1 (Cai et al., 2025), and Gemini-2.0-flash-image (Google, ), using 1,000 prompts sampled from the COYO/GBC datasets. Notably, Gemini-2.0-flash-image can perform its own prompt optimization, making it a strong closed-source baseline. As shown in Table 3, TIPO generally improves image quality and alignment relative to the baselines. For example, Lumina-2 and HiDream-I1 achieve higher aesthetic scores and lower corruption rates after TIPO optimization. Despite Gemini’s integrated optimization, TIPO still yields measurable improvements, highlighting the practical value of our lightweight, task-specific strategy. Overall, these results demonstrate that TIPO maintains robust compatibility with heterogeneous T2I models, even when their training data are unavailable and potentially mismatched with TIPO’s optimization corpus.
Efficiency
A key concern is whether TIPO’s iterative pre-sampling strategy introduces noticeable latency. Hence, we benchmark prompt-generation latency in Table 4, where TIPO reduces per-prompt latency with an improvement up to 29.4%. Details of training and inference are in Appendix H.4.
Table 4: Prompt generation latency and relative speedup
TIPO Promptist PromptExtend MagicPrompt
Avg. Time (s) 1.03 1.46 1.38 1.14
TIPO speedup vs. each +29.4% +25.6% +9.6%
Human Preference Evaluation
@@
Quantitative metrics may not fully align with human preference. Therefore, we conducted a user study based on pairwise image comparisons between the original prompt, MagicPrompt, Promptist, and TIPO on over 1,400 images, gathering preferences from 221 volunteers. As illustrated in Figure 5(a), TIPO achieved the highest overall win rate at 51.3%, significantly outperforming competitors. In out-of-domain scenarios, TIPO’s win rate increased to 52.5%, demonstrating consistently strong user preference across different contexts. For further results and statistics, please refer to Appendix G.
(a) Pairwise comparison results showing win-tie-lose percentages for overall user preference. Evaluations cover ‘All Data’ and ‘Out-of-Domain’ sets for TIPO and baseline methods.
(b) ELO ratings comparing TIPO and baseline methods across four criteria: Prompt Adherence, Image Quality, Aesthetic, and Overall. The dashed line indicates the initial ELO rating.
Figure 5: Human preference evaluation demonstrates that TIPO consistently achieves higher user preference compared to baseline prompt optimization methods (Promptist, prompttext, magicprompt, and original prompts). All evaluated images were generated using SD-3.5-Medium.
Prompt Distribution Alignment
While TIPO consistently achieves the best or competitive results in image quality, it remains unclear whether such gains arise from a closer distributional alignment between its optimized prompts and the T2I models’ training text corpora. To verify this hypothesis, we sample from two representative T2I training sets, obtaining 1,000 natural-language captions from COYO and 1,000 tag-based captions from Danbooru2023. We then encode ground-truth captions and the outputs of all compared optimization methods using two widely adopted text encoders—T5-XXL (Raffel et al., 2020a) (used in SD3, Flux, and PixArt) and jina-embeddings-v3 (Sturua et al., 2024) (a popular recent text-embedding model). Finally, we measure the embedding-space alignment between the ground-truth captions and optimized prompts with Fréchet Distance (FD) and Maximum Mean Discrepancy (MMD with RBF kernel). As shown in Table 5, baseline methods show varying extents of alignment across different prompt types, text encoders, and metrics, while TIPO maintains consistently better alignment under all settings. This indicates that TIPO aligns well with the T2I training-text distribution in a prompt-compatible and encoder-insensitive manner.
Table 5: Embedding-space distances (FD and MMD) between optimized prompts and T2I training corpora. Lower is better. Best results are in bold and second-best are underlined.
Prompt Encoder Metric TIPO Promptist MagicPrompt GPT-4o-mini
NL-short Jina FD 0.0322 0.1003 0.0385 0.1064
MMD 0.0320 0.1501 0.0624 0.1699
T5 FD 0.0704 0.2072 0.1441 0.1252
MMD 0.1438 0.2914 0.1972 0.2297
NL-trunc Jina FD 0.0309 0.1192 0.0493 0.0963
MMD 0.0359 0.1700 0.0772 0.1642
T5 FD 0.0674 0.2312 0.1884 0.1276
MMD 0.1404 0.3147 0.2270 0.2323
Tag-based Jina FD 0.1094 0.1891 0.1958 0.2479
MMD 0.1539 0.2473 0.2415 0.3050
T5 FD 0.0524 0.2080 0.2578 0.0728
MMD 0.1846 0.3573 0.3948 0.2194
Prompt–Image Alignment
To further assess the semantic consistency between optimized prompts and their generated images, we compute CLIPScore between each prompt and its corresponding SD1.5-generated image using openai/clip-vit-large-patch14-336, on the same prompt sets as in the Prompt Distribution Alignment experiment. As shown in Table 6, TIPO achieves the strongest alignment for tag-based prompts. While MagicPrompt shows clear advantages on NL-based prompts, this is likely due to its 1.8M+ SD1.5 community training corpus, where prompts are typically explicit and stylistically strong, effectively eliciting the CLIP encoder to produce high alignment scores. However, such stylistic bias is often less preferred by mainstream users. In contrast, TIPO attains competitive performance on NL-based prompts without such bias, as reflected by the win rate in Table 11, where MagicPrompt vs. TIPO = 11:48.
Table 6: CLIPScore between optimized prompts and T2I generated images. Higher is better. Best results are in bold and second-best are underlined.
Prompt Type TIPO MagicPrompt GPT-4o-mini Promptist
Tag-based 0.2217 0.1782 0.1774 0.1642
NL-short 0.2413 0.2834 0.2378 0.2347
NL-trunc 0.2310 0.2517 0.2275 0.2063
@@ 6 Conclusion
We introduced TIPO, a lightweight prompt pre-sampling framework designed for efficient real-world Text-to-Image (T2I) applications. By aligning user prompts with the intrinsic distributions of T2I training datasets, TIPO enhances semantic coherence, image fidelity, and diversity with minimal inference overhead. Experimental results show that TIPO consistently outperforms existing prompt optimization methods across multiple evaluation metrics, while extensive user studies confirm its strong alignment with human preferences. Despite these promising results, several aspects remain open for future exploration, such as generalization to out-of-distribution user input, model-specific adaptation, personalization, image-feedback-aware refinement, and large-scale model scaling. We discuss these limitations and potential extensions in Appendix J. To encourage wider adoption and facilitate reproducibility, we release our trained models and source code. We hope TIPO will inspire further advancements in efficient, scalable, and robust generative frameworks for creative systems.
Ethics Statement
Our work democratizes prompt optimization for text-to-image models, but automatic prompt enrichment can inherit biases from training data or be misused to produce misleading or harmful content. We therefore emphasize responsible use, including bias mitigation, transparency, and appropriate user controls. No human subjects or sensitive personal data were involved, and we avoided including identifiers or metadata that could raise copyright- or attribution-related concerns.
Reproducibility Statement
An anonymized repository with code and configuration files is provided in the supplementary materials, enabling reproduction of all reported results. The main paper and appendix reference the exact experimental settings, hyperparameters, random seeds, and evaluation scripts; data preprocessing steps and any additional resources needed to re-create the experiments are also documented there. For hardware, we report that all experiments were conducted on NVIDIA RTX 3090 and/or A6000 GPUs, with full details described in the appendix and repository.
@@ A.1 Danbooru2023
The Danbooru2023 dataset (Yeh, 2024a; b; nyanko202, 2023) is an extensive collection of images and their corresponding tags, compiled from the Danbooru image board. This dataset includes images annotated with particular and detailed tags, providing a rich resource for training both the Text-to-Image (T2I) and Large Language Models (LLMs) involved in the TIPO framework. The dataset contains data up to image ID 7,349,999, encompassing various visual content with granular annotations. These annotations allow for creating nuanced and precise prompts, ensuring that longer, more detailed prompts can indicate subsets of shorter prompts.